OpenStudy (anonymous):

What is the nth term of the geometric sequence that has a common ratio of 6 and 24 as its third term?

OpenStudy (anonymous):

a3=a1*6^(n-1) 24=a1*6^(3-1) since it is the third term 3/2=a1 plug in a1 to the formula an=3/2*6^(n-1)